# Imaging cross-degree-of-freedom structures in biphoton spatial-polarization hyperentanglement

### v2

@@ Abstract
Hyperentanglement, which refers to entanglement across more than one degree of freedom (DoF), is a valuable resource in photonic quantum technology. Although hyperentangled states inherently facilitate access to high-dimensional Hilbert spaces, their distinguishing feature lies in the cross-DoF structures beyond independent entanglement in each DoF, which can be leveraged to enable novel and robust information processing schemes. In particular, for a spatial-polarization hyperentangled state generated from spontaneous parametric down-conversion, the resulting polarization entanglement is structured by the spatial distribution of all interacting optical fields. Characterizing such cross-DoF structures is a crucial step towards employing hyperentangled states for practical quantum applications. Here, by performing polarization state tomography on all spatially correlated photon pairs, we have, for the first time, resolved the full-field spatial-polarization structure in a 251-dimensional hyperentangled biphoton state. We observe an entire class of near-maximally polarization-entangled states with an average concurrence of , characterized by biphoton polarization phases that exhibit radial and linear gradients induced by the transverse momenta of the down-converted photons and the pump beam, respectively. Our study lays important groundwork for further exploiting the high dimensionality and cross-DoF correlations in hyperentangled states for future quantum technologies.
1 Introduction
Hyperentanglement, which refers to the simultaneous entanglement in multiple degrees of freedom (DoFs) [1, 2], could be a key enabler of robust quantum information technologies and advance the research in fundamental aspects of quantum physics. One important reason is its inherent advantage in achieving high dimensionality. The tensor product postulate of quantum theory implies that the resulting dimensionality of a state equals the product of the dimensionalities in each constituent DoF. The benefit of harnessing hyperentanglement is particularly prominent in photonic quantum systems: while the dimensionality of polarization entanglement is inherently limited to two, and experimental factors constrain that of the spatial entanglement, the presence of spatial-polarization hyperentanglement can effectively double the dimensionality from what is achievable without it. The resulting high dimensionality can thus lead to tremendous potential for quantum information processing [3, 4]: it could boost security and noise-resilience in quantum communication channels [5, 6, 7, 8, 9], enable efficient implementations of quantum computation [10, 11, 12], improve sensitivity in metrology protocols [13, 14], and enhance noise tolerance in experimental tests on foundations of quantum theory [15, 16]. In particular, hyperentangled photons have been implemented to enable superdense coding and teleportation [17, 18, 19], enhance the noise resilience of quantum illumination [20], and improve the secure key rates of quantum communication [21, 22, 23].
@@ 1 Introduction
The implications of using hyperentangled states extend beyond multiplying entanglement dimensionalities in different DoFs. For instance, it was shown that quantum correlations across spatial and polarization DoFs can enable novel and more robust information encoding schemes and demonstrate topological resilience against environmental noise [24]. The spatially varying polarization correlations in a hyperentangled state can also be harnessed to enable quantum holographic imaging [25] or engineer photonic cluster states for efficient quantum computation [26], implying the potential of harnessing the interplay between different DoFs for novel quantum information technologies. From a fundamental perspective, studies have shown that spatial modes of the pump beam can influence the polarization entanglement produced from spontaneous parametric down-conversion (SPDC) [27, 28], indicating the opportunity of directly generating desired cross-DoF structures in a hyperentangled state, or encoding phase images in the biphoton correlations [29] by shaping the pump beam.
@@
Although a hyperentangled biphoton state can be readily generated from SPDC using a paired crystal configuration [30, 1], its potential, especially that embedded in the spatial-polarization structure, remains underexplored. The first and foremost reason is that the multiplicative increase in dimensionality poses a significant challenge to the efficiency of hyperentanglement characterization. For instance, a full state tomography for a bipartite state with local dimensionality in a single DoF requires single-outcome projective measurements [31, 4, 32]. The need to address correlations across multiple DoFs simultaneously only adds to the complexity of the problem. Without comprehensively resolving the polarization entanglement produced in each spatiotemporal mode, the rich structure produced by cross-DoF interplay reduces to a mixture of distinguishable states. As a result, the full potential of hyperentanglement becomes inaccessible, limiting its application in practical quantum devices. For instance, the work by Li et al. [27] indicates that even though each spatial mode in a multimode pump beam can generate maximum polarization entanglement through SPDC, detecting the biphotons without resolving individual spatial modes can make the entanglement appear deteriorated.
@@
Recently, several studies have employed Tpx3Cam, a data-driven camera capable of time-stamping single photons [33, 34, 35], to enable rapid characterization of photonic entangled states produced from SPDC [36, 37, 38]. However, Ref. [36, 37] have focused only on the polarization DoF or the spatial DoF, respectively, while their photon pair sources are capable of producing hyperentangled states. As a result, the potential of hyperentanglement and the embedded cross-DoF structures has not been addressed. Furthermore, the entangled states produced in Ref. [36, 38] are first spatially postselected using single-mode fibers or polarization-maintaining fibers before being imaged onto the Tpx3Cam. Doing so effectively reduces the spatial dimensionality and obscures any cross-DoF structures imposed by the generation process. Understanding how different DoFs become interlinked is essential for tailoring hyperentangled states to specific applications. Therefore, a quantitative and comprehensive characterization of these cross-DoF structures is crucial to fully harness the potential of hyperentanglement.
@@
In this work, we employ Tpx3Cam to comprehensively characterize the spatial-polarization structure in a hyperentangled biphoton state produced from SPDC. We first certify spatial-polarization hyperentanglement by measuring the entanglement in each constituent DoF. Measurements of position and momentum correlations estimate a Schmidt number in the spatial DoF to be approximately 148. Spatially-resolved polarization state tomography further confirms the generation of near-maximally polarization-entangled states across the entire spatial profile of the SPDC field, with an average concurrence of , certifying a total hyperentanglement dimensionality of 251. To the best of our knowledge, this is the highest hyperentanglement dimensionality certified thus far. To illustrate the spatial-polarization structure in the hyperentangled state, we present the first complete spatial map of the polarization state generated from SPDC, revealing how polarization entanglement depends on the transverse momenta of the down-converted photons throughout the entire spatial profile of the SPDC field. Moreover, using a weakly focused pump beam, we experimentally characterize, for the first time, the dependence of polarization entanglement on the angular spectrum of the pump beam. Our results advance the fundamental understanding of cross-DoF structures in hyperentangled states, paving the way for novel applications in high-dimensional quantum information processing.
2 Experimental setup
@@
Fig. 1(a) depicts our experimental setup. An ultraviolet continuous-wave laser emits the pump beam with a center wavelength of 405 nm, a bandwidth of 2 nm, and a power of 20 mW. The pump beam is made polarized at using a polarizing beam splitter (PBS) and a half-wave plate (HWP). A 5-mm temporal compensator (TC) quartz crystal introduces a time delay between the horizontal (H-) and vertical (V-) components of the pump beam, which pre-compensates for the temporal walk-off that the two orthogonal polarization components are expected to subsequently gather inside the BBO double-crystal [39]. The BBO crystals are each 0.5-mm-thick and identically cut for type-I phase-matching, with their optical axes oriented perpendicularly to one another [30]. A -polarized photon from the pump beam can then be down-converted in either crystal with equal probability to produce a pair of H-polarized or V-polarized photons. In the experiment, we orient the double-crystal for near-collinear phase matching. The down-converted photons are separated from the pump beam using a dichroic mirror (DM), and then probabilistically split into two arms using a non-polarizing beam splitter (BS). We ignore the cases wherein both the down-converted photons exit the same port of the BS, and denote photons that are reflected as signal and those that are transmitted as idler. This beam-splitting scheme ensures that measuring correlation between the signal arm and the idler arm resolves the cross-DoF structure over the full spatial profile of the SPDC. Although the probabilistic splitting reduces the coincidence rates by , it does not affect the quality of entanglement or the measured cross-DoF structure compared to those obtained with deterministic splitting, as can be seen later in Fig. 3 and Fig. 4. In the low-gain regime, the output state in the far-field can be written in the joint spatial-polarization basis as
@@
where represents the transverse momentum of the signal (idler) photon and are complex coefficients. The quantity represents a phase difference between the orthogonal polarization components of the state, which is dependent on the transverse momentum of the signal and the idler photons [40]. In what follows, we refer to as the biphoton polarization phase.
@@
The signal and idler photons then propagate through a quarter-wave plate (QWP) and a HWP before they are either transmitted or reflected by a PBS. The photons finally pass through a bandpass filter with a center wavelength of 800 nm and full width at half maximum (FWHM) bandwidth of 40 nm before being detected by the Tpx3Cam. To characterize the entanglement in the spatial DoF, we implement two imaging schemes for realizing measurements in the position (near-field) and momentum (far-field) bases while setting the waveplates to project both the signal and idler photons into the V-polarization basis. Using the combination of mm, mm, and mm lenses, the far-field of the crystal is imaged on the Tpx3Cam, thereby enabling measurements in the momentum basis. On the other hand, using mm images the crystal output face (near-field) onto the Tpx3Cam sensor, thereby allowing for measurements in the position basis. We acquire data in the near-field and far-field for 1 minute each using the Tpx3Cam. To characterize entanglement in the polarization DoF, we additionally acquire data for 15 different combinations of orientations of HWPs and QWPs in the far-field, each for 1 minute. Combined with the VV basis data acquired in characterizing the far-field spatial correlation, these polarization measurements allow us to reconstruct the polarization density matrices corresponding to all pairs of momentum-anticorrelated pixels through quantum state tomography and characterize polarization entanglement. The total required data acquisition time for characterizing a spatial-polarization hyperentangled state is thus minutes.
@@
The imaging sensor in Tpx3Cam comprises a pixel array with a pixel pitch of 55 m. The pixels in Tpx3Cam are data-driven and individually trigger the registration of photon incidence events when the signal amplitude exceeds a predefined threshold. The camera can be single photon sensitive with the addition of an image intensifier (Photonis Cricket) and have a single photon temporal resolution of 2 ns [35]. Time-stamping for individual photons can potentially allow for more versatility in data acquisition and analysis compared to frame-based imaging devices. During data acquisition, each incident photon could hit a cluster of pixels due to being amplified by the intensifier. To correct this, we apply a centroiding algorithm that identifies the amplitude-weighed center in each cluster as the true pixel coordinate. We use the time stamp of each centroided pixel as a reference to correct the time walk within the cluster. After centroiding and time walk correction, we apply a two-pointer technique to the sorted time stamps of signal and idler photons and identify events detected within a 10-ns time window as photon coincidences [36]. Although the time walk correction algorithm can reduce the temporal resolution to 8 ns [41, 42], it should not significantly affect the efficiency of the subsequent coincidence counting with our choice of longer time window. As a result, we obtain an average coincidence rate of approximately 2.7 Hz between two spatially correlated -pixel regions, which have similar sizes with the correlation widths measured at the Tpx3Cam sensor plane (see supplemental document for details), and an overall coincidence rate of 3200 Hz between the full signal and idler fields. We note that although the experimentally measured coincidence rate is limited by the overall photon detection efficiency of the Tpx3Cam system [41], our choice of 1-minute data acquisition time has allowed us to obtain enough counts for the relevant results to be statistically valid.
3 Results and discussions
3.1 Certifying spatial-polarization hyperentanglement
We first demonstrate the presence of hyperentanglement by certifying its high dimensionality. The dimensionality of the full spatial-polarization hyperentangled biphoton state depends primarily on the Schmidt number in the spatial DoF. While past studies have estimated this number based on certain prior assumptions about the two-photon state [43, 44, 45, 46, 47], some recent studies have demonstrated assumption-free protocols for certifying high dimensionality [48, 49]. In particular, it was shown that measurements in two mutually unbiased bases (MUBs) can efficiently estimate a lower bound for the Schmidt number of an entangled state, thereby certifying high-dimensional entanglement without the need for prior assumptions [48]. In our case, we show that the entanglement dimensionality in the spatial DoF can be certified by measuring the photon correlation in the discrete position basis and the discrete momentum basis, similar to the approach implemented in Ref. [50].
We denote the spatial modes of the signal(idler) photons in the discrete momentum basis as and in the discrete position basis as . Based on the experimentally measured spatial correlation widths (see supplemental document for details), individual discrete modes in the Tpx3Cam sensor plane are better approximated by -pixel large superpixels with 2-pixel separations between adjacent superpixels. We note that since this protocol is sensitive to noise introduced by cross-talk between non-conjugate modes, a different choice in the superpixel size and separation could lead to significant underestimation of the dimensionality. As shown in Fig. 2(a) and 2(c), we select two sets of superpixels evenly distributed over the regions illuminated by signal and idler photons. The cross-correlations between spatial modes of signal and idler photons are used to calculate , a lower bound for the fidelity of the experimentally measured state to a maximally entangled state . The dimensionality of the spatial entanglement is certified to be at least if satisfies [48]
Upon analyzing the cross-correlations between selected spatial modes of signal and idler photons, we obtain the correlation matrices shown in Fig. 2(b) and 2(d). The fidelity lower bound is then calculated to be (see supplemental document for details). In other words, a maximally entangled state has to have more than 147 dimensions to be transformable into our measured state through local operations and classical communication (LOCC). Therefore, the entanglement dimensionality in the spatial DoF is certified to be at least 148.
@@ 3.1 Certifying spatial-polarization hyperentanglement
We now characterize the polarization entanglement of the biphotons produced in each spatial mode of the SPDC, thereby certifying spatial-polarization hyperentanglement. Using the setup depicted in Fig. 1(a), we perform quantum state tomography [31] using experimentally measured polarization correlation between each pair of momentum-anticorrelated (diametrically-opposite) superpixels in the far-field and reconstruct the corresponding polarization density matrix. We then compute the concurrence using , where ’s are the eigenvalues of a Hermitian matrix derived from applying Pauli-y operations on reconstructed density matrices [51]. The biphoton polarization phases are computed as the phase of the elements of the reconstructed density matrices.

### v3

# Full-field mapping of spatially varying polarization entanglement generated from spontaneous parametric down-conversion

@@ Abstract
Two-photon states generated from spontaneous parametric down-conversion (SPDC) can display entanglement in all degrees of freedom (DoFs) of light, including spatial, temporal, and polarization. The coupling between different DoFs of a two-photon state has been shown to display rich structures that enable novel and robust information processing schemes. While existing literature has studied these couplings by post-selecting the SPDC field, a comprehensive understanding of the inherent spatial-polarization coupling produced in the SPDC process is still lacking. This work produces a full spatial map of the polarization entanglement generated across the entire SPDC field. We observe an entire class of near-maximally polarization-entangled states with an average concurrence of , which, together with a certified spatial dimensionality of 148, could potentially offer access to a 251-dimensional hyperentanglement. The spatial-polarization coupling manifests as radially or linearly varying polarization-entangled states, whose wavefunctions are dependent on the transverse momenta of the down-converted photons and the pump beam, respectively. Our study lays important groundwork for further exploiting the coupling between entanglement in different DoFs for future quantum technologies.
I Introduction
Spontaneous parametric down-conversion (SPDC) has been a major workhorse for generating entangled tw-photon states [20, 5, 45], enabling numerous applications in quantum information processing. In SPDC, a photon from a pump beam, which has a higher optical frequency, interacts with a second-order nonlinear medium to be annihilated and produce a pair of down-converted photons with lower frequencies [5]. The down-converted photons, which are commonly referred to as signal and idler, can exhibit entanglement in all degrees of freedom (DoFs) of light, including spatial, temporal, and polarization.
@@ I Introduction
Studying the coupling between different DoFs is a crucial step towards a comprehensive understanding of SPDC-based entanglement generation and engineering novel entangled photon states for practical applications [21, 34]. An important aspect of this subject is the coupling between spatial and polarization DoFs in SPDC. Such cross-DoF couplings can be exploited to enable novel and efficient quantum information applications. For instance, it was shown that quantum correlations across spatial and polarization DoFs can enable novel and more robust information encoding schemes and demonstrate topological resilience against environmental noise [40]. Moreover, the spatially varying polarization correlations in a photonic entangled state can also be harnessed to enable quantum holographic imaging [9] or engineer photonic cluster states for efficient quantum computation [46, 6]. From a fundamental perspective, studies have shown that spatial modes of the pump beam can influence the polarization entanglement of the down-converted photons [30, 53], indicating the opportunity of directly engineering the couplings between different DoFs in a photonic entangled state, or encoding phase images in the two-photon correlations [47] by shaping the pump beam.
@@
The spatial-polarization coupling in SPDC commonly manifests as two-photon states with spatially varying polarization entanglement. Although one can readily produce these features from SPDC using a paired crystal configuration [27, 2], the large spatial dimensionality of the SPDC field hinders the efficient characterization of them. For instance, a full state tomography for a bipartite state with local dimensionality in a single DoF requires single-outcome projective measurements [24, 16, 22]. Since the total dimensionality of a quantum state scales with the product of dimensionality in each constituent DoF, the need to address correlations, both within and between different DoFs, only adds to the complexity of the problem. Without comprehensively resolving the polarization entanglement produced in each spatial mode, the rich structures resulting from the coupling between different DoFs reduce to a mixture of distinguishable states. As a result, the full potential of SPDC-induced spatial-polarization coupling becomes inaccessible, limiting its application in practical quantum devices. For instance, a recent work indicates that even though each spatial mode in a multimode pump beam can generate maximum polarization entanglement through SPDC, detecting the two-photon states without resolving individual spatial modes can make the entanglement appear deteriorated [30].
@@
Over the past decade, several studies have employed Tpx3Cam, a data-driven camera capable of time-stamping single photons [15, 39, 38], to enable rapid characterization of photonic entangled states produced from SPDC [23, 7, 18]. However, these works have not fully explored the coupling between different DoFs. Specifically, Ref. [23, 7] have focused only on the polarization DoF or the spatial DoF, respectively, while their photon pair sources are capable of producing spatially varying polarization-entangled states. As a result, the potential of the embedded spatial-polarization coupling has not been addressed. Furthermore, the entangled states produced in Ref. [23, 18] are first spatially postselected using single-mode fibers or polarization-maintaining fibers before being imaged onto the Tpx3Cam, and any spatial structures in the resulting polarization are imprinted by external modulation devices. Doing so effectively reduces the spatial dimensionality and obscures any spatial-polarization coupling imposed by the generation process.
@@
In this work, we employ Tpx3Cam to comprehensively characterize the spatially varying polarization entanglement that is natively produced from SPDC. In contrast to earlier works, we image the full spatial field of SPDC without any post-processing on the resulting two-photon states. By performing spatially-resolved polarization state tomography, we confirm the generation of near-maximally polarization-entangled states across the entire spatial profile of the SPDC field, with an average concurrence of . Importantly, we present a complete spatial map of the polarization state generated from SPDC, revealing how polarization entanglement depends on the transverse momenta of the down-converted photons throughout the entire spatial profile of the SPDC field. Moreover, using a weakly focused pump beam, we experimentally characterize the dependence of polarization entanglement on the angular spectrum of the pump beam. Additionally, we have estimated a maximum attainable dimensionality of spatial entanglement to be 148, which, combined with the effective dimensionality in the polarization DoF, indicates a total attainable dimensionality of 251. Our results advance the fundamental understanding of the interplay between different DoFs of a photonic entangled state, paving the way for novel applications in photonic quantum information processing.
II Experimental setup
@@
Fig. 1(a) depicts our experimental setup. An ultraviolet continuous-wave laser emits the pump beam with a center wavelength of 405 nm, a bandwidth of 2 nm, and a power of 20 mW. The pump beam is made polarized at using a polarizing beam splitter (PBS) and a half-wave plate (HWP). A 5-mm temporal compensator (TC) quartz crystal introduces a time delay between the horizontal (H-) and vertical (V-) components of the pump beam, which pre-compensates for the temporal walk-off that the two orthogonal polarization components are expected to subsequently gather inside the BBO double-crystal [33]. The BBO crystals are each 0.5-mm-thick and identically cut for type-I phase-matching, with their optical axes oriented perpendicularly to one another [27]. A -polarized photon from the pump beam can then be down-converted in either crystal with equal probability to produce a pair of H-polarized or V-polarized photons. In the experiment, we orient the double-crystal for near-collinear phase matching. The down-converted photons are separated from the pump beam using a dichroic mirror (DM), and then probabilistically split into two arms using a non-polarizing beam splitter (BS). We ignore the cases wherein both the down-converted photons exit the same port of the BS, and denote photons that are reflected as signal and those that are transmitted as idler. This beam-splitting scheme ensures that measuring the correlation between the signal arm and the idler arm resolves the full spatial profile of the SPDC. Although the probabilistic splitting reduces the coincidence rates by , it does not affect the quality of entanglement or the measured spatial-polarization structure compared to those obtained with deterministic splitting, as can be seen later in Fig. 2 and Fig. 4. In the low-gain regime, the output state in the far-field can be written in the joint spatial-polarization basis as
@@
where represents the transverse momentum of the signal (idler) photon and are complex coefficients. The quantity represents a phase difference between the orthogonal polarization components of the state, which is dependent on the transverse momentum of the signal and the idler photons [1]. In what follows, we refer to as the two-photon polarization phase.
@@
The signal and idler photons then propagate through a quarter-wave plate (QWP) and a HWP before they are either transmitted or reflected by a PBS. The photons finally pass through a bandpass filter with a center wavelength of 800 nm and full width at half maximum (FWHM) bandwidth of 40 nm before being detected by the Tpx3Cam. To characterize entanglement in the polarization DoF, we measure the polarization correlation between each pair of momentum-correlated signal and idler photons. We first image the far-field of the crystal onto the Tpx3Cam using the combination of mm, mm, and mm lenses, and then acquire data for 16 different combinations of orientations of HWPs and QWPs, each for 1 minute. These polarization measurements allow us to reconstruct the polarization density matrices corresponding to all pairs of momentum-anticorrelated photons through quantum state tomography, which characterizes polarization entanglement. To characterize the state dimensionality in the spatial DoF, we measure the spatial correlation of signal and idler photons in the position basis. We do so by replacing the mm lenses with mm ones, thereby imaging the crystal output face (near-field) onto the Tpx3Cam sensor. We then set the HWPs and QWPs to project the polarization in both arms into V and acquire data for 1 minute. Combined with the 16 measurements acquired in polarization characterization, the total required data acquisition time is thus minutes.
@@
The imaging sensor in Tpx3Cam comprises a pixel array with a pixel pitch of 55 m. The pixels in Tpx3Cam are data-driven and individually trigger the registration of photon incidence events when the signal amplitude exceeds a predefined threshold. The camera can be single-photon sensitive with the addition of an image intensifier (Photonis Cricket) and has a single photon temporal resolution of 2 ns [38]. Time-stamping for individual photons can potentially allow for more versatility in data acquisition and analysis compared to frame-based imaging devices. During data acquisition, each incident photon could hit a cluster of pixels due to being amplified by the intensifier. To correct this, we apply a centroiding algorithm that identifies the amplitude-weighed center in each cluster as the true pixel coordinate. We use the time stamp of each centroided pixel as a reference to correct the time walk within the cluster. After centroiding and time walk correction, we apply a two-pointer technique to the sorted time stamps of signal and idler photons and identify events detected within a 10-ns time window as photon coincidences [23]. Although the time walk correction algorithm can reduce the temporal resolution to 8 ns [48, 54], it should not significantly affect the efficiency of the subsequent coincidence counting with our choice of a longer time window. As a result, we obtain an average coincidence rate of approximately 2.7 Hz between two spatially correlated -pixel regions, which have similar sizes with the correlation widths measured at the Tpx3Cam sensor plane (see Appendix A for details). Throughout this paper, we refer to each -pixel region as a superpixel. Moreover, we observe an overall coincidence rate of 3200 Hz between the full signal and idler fields. We note that although the experimentally measured coincidence rate is limited by the overall photon detection efficiency of the Tpx3Cam system [48], our choice of 1-minute data acquisition time has allowed us to obtain enough counts for the relevant results to be statistically valid.
III Results and discussions
III.1 Full spatial mapping of polarization entanglement
@@
Using the setup depicted in Fig. 1(a), we perform quantum state tomography [24] using experimentally measured polarization correlation between each pair of momentum-anticorrelated (diametrically-opposite) superpixels in the far-field and reconstruct the corresponding polarization density matrix. We then compute the concurrence using , where ’s are the eigenvalues of a Hermitian matrix derived from applying Pauli-y operations on reconstructed density matrices [49]. The two-photon polarization phases are computed as the phase of the elements of the reconstructed density matrices.
Fig. 2 depicts the full spatial maps of the concurrence and two-photon polarization phase in the far field. For illustration purposes, we have mapped all signal-idler pairs to the central coordinates of the superpixels of signal photons. In Fig. 2(c), we also display a two-photon polarization density matrix, which is reconstructed at the center of the SPDC field, with the corresponding signal-idler pair indicated by a black square in Fig. 2(a-b). We calculate an average concurrence of 0.8303 0.0004, which implies strong polarization entanglement with a spatially varying structure across the entire SPDC field.
It is important to note that in Fig. 2(b), the two-photon polarization phase is highly uniform along the azimuthal direction while displaying a gradient along the radial direction, indicating highly indistinguishable polarization entanglement at each circular region concentric with the SPDC field. Consequently, if each corresponding spatial mode is also highly indistinguishable from one another, the two-photon state can be seen as being hyperentangled [28, 2], which is a tensor product of (i) spatially entangled discrete momentum modes occupying the full spatial field of SPDC; and (ii) polarization-entangled states with a well-defined two-photon polarization phase . In the following discussions, we characterize the spatial distinguishability between all correlated spatial modes by certifying the spatial dimensionality of the two-photon state, thereby estimating an attainable total dimensionality of the full SPDC field in the joint spatial-polarization basis.
III.2 Certifying total attainable dimensionality of in the joint spatial-polarization bases
We characterize the spatial dimensionality of a two-photon state using its Schmidt number. While past studies have estimated this number based on certain prior assumptions about the two-photon state [29, 11, 12, 10, 17], some recent studies have demonstrated assumption-free protocols for certifying high dimensionality [3, 14]. In particular, it was shown that measurements in two mutually unbiased bases (MUBs) can efficiently estimate a lower bound for the Schmidt number of an entangled state, thereby certifying high-dimensional entanglement without the need for prior assumptions [3]. In our case, we show that the entanglement dimensionality in the spatial DoF can be certified by measuring the photon correlation in the discrete position basis and the discrete momentum basis, similar to the approach implemented in Ref. [35].
We denote the spatial modes of the signal(idler) photons in the discrete momentum basis as and in the discrete position basis as . Based on the experimentally measured spatial correlation widths (see Appendix B for details), individual discrete modes in the Tpx3Cam sensor plane are better approximated by -pixel large superpixels with 2-pixel separations between adjacent superpixels. We note that since this protocol is sensitive to noise introduced by cross-talk between non-conjugate modes, a different choice in the superpixel size and separation could lead to significant underestimation of the dimensionality. As shown in Fig. 3(a-b), we select two sets of superpixels evenly distributed over the regions illuminated by signal and idler photons. The cross-correlations between spatial modes of signal and idler photons are used to calculate , a lower bound for the fidelity of the experimentally measured state to a maximally entangled state . The dimensionality of the spatial entanglement is certified to be at least if satisfies [3]
Upon analyzing the cross-correlations between selected spatial modes of signal and idler photons, we obtain the correlation matrices shown in Fig. 3(c-d). The fidelity lower bound is then calculated to be (see Appendix B for details). In other words, a maximally entangled state has to have more than 147 dimensions to be transformable into our measured state through local operations and classical communication (LOCC). Therefore, the entanglement dimensionality in the spatial DoF is certified to be at least 148, indicating high indistinguishability between all correlated spatial modes.
In principle, the polarization phase in Eq. 1 can be compensated using spatial light modulators (SLMs) so that it becomes uniform also across all radial positions without reducing the polarization entanglement [9]. Consequently, we expect the attainable total dimensionality of our state to be equal to that of a global hyperentangled state, which is a product of the average dimensionality in the polarization DoF and the certified dimensionality in the spatial DoF, as underscored by the tensor product postulate of quantum physics.